Two wires A and B have same lengths and material , have their cross sectional areas 1:4 , what would be the ratio of heat produced in these wires when the voltage across each is constant.
1
Expert's answer
2014-03-15T03:45:50-0400
Q1 = (U1*U1)*t/R1 R1 = Ro1 * L1/S1
Q2 = (U2*U2)*t/R2 R2 = Ro2 * L2/S2 S2=4* S1
Q1/Q2 =R2/R1 = S1/S2 = 1/4